OpenStudy (anonymous):

simplify (-Sin^2(x)+sin(x))/(1+sin(x))

OpenStudy (mathteacher1729):

Is there anything you can factor out of the numerator? \(\sin(x)-\sin(x)*\sin(x)\) ?

OpenStudy (anonymous):

sinx(sinx+1)/(sinx+1)=sinx

OpenStudy (anonymous):

the original problem is (Sin^2(-x)-sin(-x))/(1-sin(-x))

OpenStudy (mathteacher1729):

Sin(-x) = -Sin(x) so your original question is equivalent. :)

OpenStudy (anonymous):

well yeah - but see gave an answer that disappeared a negative @_@.. I have gotten the equation to the point of looking like this: -sin^2(x)sin(x)-sin^2(x)..Any ideas?